OpenStudy (anonymous):

How Do You Write 1/8 As A Percent ? @dmezzullo

OpenStudy (anonymous):

.125=12.5%

OpenStudy (perl):

1/8 x 100

OpenStudy (anonymous):

1/8 can be thought of as 1 divided by 8. Think of the fraction bar as a division symbol in which you divide the numerator by the denominator. 1/8 = 0.125 . Example: 8/1 = 8 divided by 1 = 8 Example: 3/36 = 3 divided by 36 = .08333 Example: 36/3 = 36 divided by 3 = 12 . Now to change any of these decimals into a percentage, just multiply by 100%. 1/8 = 0.125 X 100% = 12.5% 3/36 = .0833 x 100% = 8.33% 36/3 = 12 x 100% =1200%

OpenStudy (perl):

1/8 x 100 =12.5 %

OpenStudy (anonymous):

0.8 ! iGuess.. lol

OpenStudy (anonymous):

You multiply what you have by 100/100, and it gives you \[\frac{1}{8}=\frac{100}{8}\frac{1}{100}=\frac{100}{8} \%=12.5 \%\]

OpenStudy (unklerhaukus):

1=100% 1/2=100%/2=50% 1/4=50%/2=25% 1/8=25%/2=

OpenStudy (perl):

% literally means divide by 100 , so 12.5% = 12.5 / 100 = .125 which is 1/8

OpenStudy (anonymous):

Hum, interesting line of though @UnkleRhaukus

OpenStudy (perl):

since you want your answer to be a percent, you have to multiply by 100 12.5% = 12.5 /100 = .125